OpenStudy (anonymous):

how to find positive root of sin^-1(x) + cos^-1(4x)= pi/6

OpenStudy (anonymous):

hmm i wonder cant think of anything other than maybe taking the sine of both sides? maybe?

OpenStudy (anonymous):

satellite73 u meant that by multiplying sine to both side....??

OpenStudy (anonymous):

i think that is what you are going to have to do \[\sin(\arcsin(x)+\arccos(4x))=\frac{1}{2}\] as a first step

OpenStudy (anonymous):

the expression on the left can be written without trig functions via \[\sin(\alpha+\beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\]

OpenStudy (anonymous):

so arcsine(x) we assume as ALPHA n arccos(4X) as BETA is'nt

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

@mazfar24 were you able to get solution?

OpenStudy (anonymous):

no...:1

OpenStudy (anonymous):

i got stuck when it become like this \[\sin \alpha \cos \beta +\cos \alpha \sin \beta =1/2\]

OpenStudy (freckles):

don't forgot what alpha and beta are

OpenStudy (freckles):

\[\sin(\arcsin(x))\cos(\arccos(4x))+\cos(\arcsin(x))\sin(\arccos(4x))=\frac{1}{2}\]

OpenStudy (freckles):

what is sin(arcsin(x))=?

OpenStudy (anonymous):

what does u meant...erm, i think sin(arcsin(X)) = to x because i tried to using phatogoras theorem...am i correct

OpenStudy (freckles):

yes that is equal to x what is cos(arccos(4x))=?

OpenStudy (anonymous):

i guest it is = 4x, am i correct??

OpenStudy (freckles):

ok now you need to evaluate cos(arcsin(x)) and also evaluate sin(arccos(4x)) do you know how to do that?

OpenStudy (dumbcow):

\[\cos(\arcsin x) = \sqrt{1-\sin^2 (\arcsin x)} = \sqrt{1-x^2}\]

OpenStudy (anonymous):

i think my answer is same like dumbcow...\[\sqrt{1-x ^{2}}\]

OpenStudy (dumbcow):

this is from the pythagorean identity: sin^2 +cos^2 = 1 so similarly \[\sin(\arccos 4x) = \sqrt{1-\cos^2 (\arccos 4x)} = \sqrt{1-16x^2}\]

OpenStudy (anonymous):

am i correct

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

but y ur answer is \[\sqrt{1-16x^{2}}\]

OpenStudy (dumbcow):

because you are squaring the "4x"

OpenStudy (anonymous):

okey, that answer is for sin(arccos4x) is'nt...so then what we must do

OpenStudy (dumbcow):

yes it is, my post clearly states that ok so now we have: \[4x^2 + \sqrt{1-x^2} \sqrt{1-16x^2} = \frac{1}{2}\] isolate sqrt term \[\sqrt{1-x^2} \sqrt{1-16x^2} = \frac{1}{2} - 4x^2\] square both sides to get rid of sqrt \[(1-x^2)(1-16x^2) = ( \frac{1}{2} - 4x^2)^2\] solve like you would a quadratic equation

OpenStudy (anonymous):

am i will get this answer \[4+\sqrt{67}/34\] n [4-\sqrt{67}/34\]

OpenStudy (dumbcow):

hmm thats not what i get what do get after distributing and combining like terms?

OpenStudy (anonymous):

i got \[17x^{2}-4x-3/4=0\]

OpenStudy (anonymous):

am i correct??

OpenStudy (dumbcow):

ahh not quite, that should be "4x^2" not just 4x

OpenStudy (anonymous):

y...can u show how u get that..please...

OpenStudy (dumbcow):

\[\rightarrow 1 -x^2 -16x^2 +16x^4 = \frac{1}{4} -4x^2 +16x^4\] \[1-17x^2 = \frac{1}{4} -4x^2\] \[13x^2 = \frac{3}{4}\]